# Towards Robust Pseudo-Label Learning in Semantic Segmentation: An Encoding Perspective

### v1

@@ Appendix C Analysis of Coding Strategy
To further study the different selections for coding strategy, we conduct experiments built with denoted as $M_{mmd}$ and $M_{text}$ respectively, and evaluate performance on UDA setting with DAFormer [hoyer2022daformer] and SSL setting with UniMatch [yang2023revisiting]. The results are shown in Table 5 and Table 6. We also We present a comprehensive comparison in Figure 6. Each code design has its own benefits:
@@
Table 5: Results on Cityscapes of UDA setting built with DAFormer [hoyer2022daformer].
@@
Table 6: Results on Pascal of SSL setting bulit with UniMatch [yang2023revisiting].
@@ Appendix G More Experiments on SSL
Challenging Benchmarks. Follow the basic setting of UniMatch [yang2023revisiting], the initial learning rate is set as 0.005 and 0.004 for Cityscapes and COCO respectively, with a SGD optimizer. The model is trained for 240, and 30 epochs under a poly learning rate scheduler. The training resolution is set as 801, and 513 for these two datasets. We adopt the Xception-65 [chollet2017xception] as backbone when trained on COCO. In Table 8, we implement the experiments on Cityscapes. ECOCSeg consistently outperforms baselines with gains ranging from 1.0% to 3.2%, especially on most challenging 1/16 partition, verifying its robustness and generalization ability. We also show the results on COCO in Table 9, where ECOCSeg outperforms the baseline with gains ranging from 1.7% to 2.6%. We find ECOC is efficient to handle a larger number of classes due to its binary decoupling property. Theoretically, a length of $K=\log_{2}N$ is sufficient to represent classes, while excessively long encodings can lead to redundancy. Empirically, the optimal code length is roughly $K=10\log_{2}N$ [allwein2000reducing]. Based on this empirical rule, we use $K=60$ for COCO (81 classes) and conduct experiments with UniMatch, observing significant improvements.
Table 8: SSL segmentation performance on Cityscapes. The 321321 and $513$ denote the training resolution.
@@
U2PL [wang2022semi] 70.6 73.0 76.3 77.2 74.9 76.5 78.5 79.1
AugSeg [zhao2023augmentation] 73.7 76.4 78.7 79.3 75.2 77.8 79.5 80.4
DAW [sun2023daw] 75.2 77.5 79.1 79.5 76.6 78.4 79.8 80.6
RankMatch [mai2024rankmatch] 75.4 77.7 79.2 79.5 77.1 78.6 80.0 80.7
Fixmatch [sohn2020fixmatch] 72.6 75.7 76.8 78.2 74.2 76.2 77.2 78.4
@@
PseudoSeg [zoupseudoseg] 29.8 37.1 39.1 41.8 43.6
PC2Seg [zhong2021pixel] 29.9 37.5 40.1 43.7 46.1
AllSpark [wang2024allspark] 34.1 41.6 45.4 49.5 50.9
UniMatch [yang2023revisiting] 31.9 38.9 44.4 48.2 49.8
@@
Powerful Baselines. To demonstrate the compatibility of our method with modern architectures and training paradigms, we conduct additional experiments on UniMatch V2 [yang2025unimatch] using the Pascal VOC high-quality set, with a DINOv2-S [oquab2023dinov2] encoder. The results are summarized in Table 10. These results confirm that ECOCSeg consistently improves performance, even when built upon strong semi-supervised learning baselines with large-scale pre-training.
@@
AugSeg [zhao2023augmentation] (RN-101) 71.1 75.5 78.8 80.3 81.4
CorrMatch [sun2024corrmatch] (RN-101) 76.4 78.5 79.4 80.6 81.8
BeyondPixels [howlader2024beyond] (RN-101) 77.3 78.6 79.8 80.8 81.7
UniMatch V2 [yang2025unimatch] (DINOv2-S) 79.0 85.5 85.9 86.7 87.8
@@
Remote Sensing: We integrate ECOCSeg into UniMatch [yang2023revisiting] (PSPNet [zhao2017pyramid]) for binary change detection on the WHU-CD dataset [ji2018fully]. To thoroughly evaluate the effectiveness of ECOCSeg, we split the WHU-CD dataset into three subsets following previous methods [yang2023revisiting]: a training set containing 5,947 images,a verification setwith 743 images,and a test set comprising 744 images. The results are shown in Table 11.
@@
S4GAN [mittal2019semi] 18.3 62.2 70.8 76.4
SemiCDNet [peng2020semicdnet] 51.7 62.0 66.7 75.9
SemiCD [bandara2022revisiting] 65.8 68.1 74.8 77.2
UniMatch [yang2023revisiting] 77.5 78.9 82.9 84.4
@@
Medical Imaging: We also evaluate ECOCSeg on the ACDC dataset [bernard2018deep], a four-class cardiac MRI segmentation task. Results using UniMatch with a UNet [ronneberger2015u] backbone are shown in Table 12.
@@
UA-MT [yu2019uncertainty] - 61.0 81.5
CPS [chen2021semi] - 60.3 83.3
CNN&Trans [luo2022semi] - 65.6 86.4
UniMatch [yang2023revisiting] 85.4 88.9 89.9
@@ Appendix H Analysis of Quality Estimate
In pseudo-label learning, we need a quality estimate $q(p_{ij})$ to control the optimization process for $\mathcal{L}^{u}$, where $p_{ij}$ is confidence and typically defined by maximum class probability. In DAFormer [hoyer2022daformer], it is defined by:
@@
where $\cdot$] denotes the Iverson bracket, and this weight is applied to all pixels of the entire image simultaneously. In UniMatch [yang2023revisiting], it is defined through a threshold filtering way:
@@ Appendix I Influence of Parameters Setting
Figure 8: Reliability diagrams for DeepLabv3+ [chen2018encoder] based on one-hot encoding (a) and ECOCSeg (b) on Pascal val.
In this section, we further study the model calibration of ECOCSeg and compare it with the one-hot encoding form. We implement DeepLabv3+ [chen2018encoder] based on these two paradigms and evaluate on Pascal val. In Fig. 8, we present the reliability diagrams, which plot the expected pixel accuracy as a the function of confidence, and calculate the Expected Calibration Error (ECE) [guo2017calibration]:
@@
where we divide the bins with an interval of 0.1. Note that the reliability diagram is obtained through the bit-wise way in ECOCSeg, meaning that every pixel will provide $K$ samples in total. As observed, ECOCSeg demonstrates smaller gaps between the expected accuracy and confidence, indicating superior calibration of predictions. While the one-hot encoding form is typically notorious for inflating the probability of the predicted class [sensoy2018evidential] and suffers higher calibration error accordingly, ECOCSeg exhibits better reliability and interpretability through fine-grained bit-level label representation.
@@
$[1]$ [chang2020weakly] Weakly-supervised Discriminative regions yield incomplete pseudo-labels Alternating optimization: 1) cluster sub-classes per class; 2) train a sub-class classifier Sub-class discovery at the feature level
$[2]$ [jiang2022prototypical] Domain adaptive Prior works focus on intra-class alignment without explicitly modeling inter-class structure Online prototype update for contrastive learning Structure modeling via class prototypes
$[3]$ [zhang2021prototypical] Domain adaptive Improve pseudo-labels via feature clustering Prototype-based denoising Feature-space method
$[4]$ [shen2023diga] Domain adaptive Avoid manual thresholds Symmetric distillation & consensus Threshold-free, feature-level
$[5]$ [zhang2021multiple] Domain adaptive Fuse online-offline pseudo-labels Unified multi-branch fusion Fusion in feature space
$[6]$ [ning2021multi] Active Domain adaptive Using a small amount of labeled target data to guide adaptation Anchor-based soft alignment Feature-space alignment
$[7]$ [mendel2020semi] Semi-supervised Treat prediction error as a learnable correction term Introduce a correction network that learns to refine predictions with residual errors. Prediction refinement via residual correction
$[8]$ [du2022learning] Semi-supervised Reduce confirmation bias Improve pseudo-labels by peeking at future model states Modification of the teacher model
@@
Figure 10: More qualitative comparison built with UniMatch [yang2023revisiting] on SSL benchmark of Pascal. The significant improvements are marked with dotted boxes.

### v2

@@ Appendix C Analysis of Coding Strategy
To further study the different selections for coding strategy, we conduct experiments built with denoted as $M_{mmd}$ and $M_{text}$ respectively, and evaluate performance on UDA setting with DAFormer [34] and SSL setting with UniMatch [93]. The results are shown in Table 5 and Table 6. We also We present a comprehensive comparison in Figure 6. Each code design has its own benefits:
@@
Table 5: Results on Cityscapes of UDA setting built with DAFormer [34].
@@
Table 6: Results on Pascal of SSL setting bulit with UniMatch [93].
@@ Appendix G More Experiments on SSL
Challenging Benchmarks. Follow the basic setting of UniMatch [93], the initial learning rate is set as 0.005 and 0.004 for Cityscapes and COCO respectively, with a SGD optimizer. The model is trained for 240, and 30 epochs under a poly learning rate scheduler. The training resolution is set as 801, and 513 for these two datasets. We adopt the Xception-65 [18] as backbone when trained on COCO. In Table 8, we implement the experiments on Cityscapes. ECOCSeg consistently outperforms baselines with gains ranging from 1.0% to 3.2%, especially on most challenging 1/16 partition, verifying its robustness and generalization ability. We also show the results on COCO in Table 9, where ECOCSeg outperforms the baseline with gains ranging from 1.7% to 2.6%. We find ECOC is efficient to handle a larger number of classes due to its binary decoupling property. Theoretically, a length of $K=\log_{2}N$ is sufficient to represent classes, while excessively long encodings can lead to redundancy. Empirically, the optimal code length is roughly $K=10\log_{2}N$ [2]. Based on this empirical rule, we use $K=60$ for COCO (81 classes) and conduct experiments with UniMatch, observing significant improvements.
Table 8: SSL segmentation performance on Cityscapes. The $321$ and $513$ denote the training resolution.
@@
U2PL [87] 70.6 73.0 76.3 77.2 74.9 76.5 78.5 79.1
AugSeg [102] 73.7 76.4 78.7 79.3 75.2 77.8 79.5 80.4
DAW [77] 75.2 77.5 79.1 79.5 76.6 78.4 79.8 80.6
RankMatch [58] 75.4 77.7 79.2 79.5 77.1 78.6 80.0 80.7
Fixmatch [72] 72.6 75.7 76.8 78.2 74.2 76.2 77.2 78.4
@@
PseudoSeg [106] 29.8 37.1 39.1 41.8 43.6
PC2Seg [104] 29.9 37.5 40.1 43.7 46.1
AllSpark [82] 34.1 41.6 45.4 49.5 50.9
UniMatch [93] 31.9 38.9 44.4 48.2 49.8
@@
Powerful Baselines. To demonstrate the compatibility of our method with modern architectures and training paradigms, we conduct additional experiments on UniMatch V2 [94] using the Pascal VOC high-quality set, with a DINOv2-S [63] encoder. The results are summarized in Table 10. These results confirm that ECOCSeg consistently improves performance, even when built upon strong semi-supervised learning baselines with large-scale pre-training.
@@
AugSeg [102] (RN-101) 71.1 75.5 78.8 80.3 81.4
CorrMatch [73] (RN-101) 76.4 78.5 79.4 80.6 81.8
BeyondPixels [33] (RN-101) 77.3 78.6 79.8 80.8 81.7
UniMatch V2 [94] (DINOv2-S) 79.0 85.5 85.9 86.7 87.8
@@
Remote Sensing: We integrate ECOCSeg into UniMatch [93] (PSPNet [101]) for binary change detection on the WHU-CD dataset [39]. To thoroughly evaluate the effectiveness of ECOCSeg, we split the WHU-CD dataset into three subsets following previous methods [93]: a training set containing 5,947 images,a verification setwith 743 images,and a test set comprising 744 images. The results are shown in Table 11.
@@
S4GAN [61] 18.3 62.2 70.8 76.4
SemiCDNet [65] 51.7 62.0 66.7 75.9
SemiCD [4] 65.8 68.1 74.8 77.2
UniMatch [93] 77.5 78.9 82.9 84.4
@@
Medical Imaging: We also evaluate ECOCSeg on the ACDC dataset [5], a four-class cardiac MRI segmentation task. Results using UniMatch with a UNet [68] backbone are shown in Table 12.
@@
UA-MT [97] - 61.0 81.5
CPS [12] - 60.3 83.3
CNN&Trans [57] - 65.6 86.4
UniMatch [93] 85.4 88.9 89.9
@@ Appendix H Analysis of Quality Estimate
In pseudo-label learning, we need a quality estimate $q(p_{ij})$ to control the optimization process for $\mathcal{L}^{u}$, where $p_{ij}$ is confidence and typically defined by maximum class probability. In DAFormer [34], it is defined by:
@@
where $\cdot$] denotes the Iverson bracket, and this weight is applied to all pixels of the entire image simultaneously. In UniMatch [93], it is defined through a threshold filtering way:
@@ Appendix I Influence of Parameters Setting
Figure 8: Reliability diagrams for DeepLabv3+ [10] based on one-hot encoding (a) and ECOCSeg (b) on Pascal val.
In this section, we further study the model calibration of ECOCSeg and compare it with the one-hot encoding form. We implement DeepLabv3+ [10] based on these two paradigms and evaluate on Pascal val. In Fig. 8, we present the reliability diagrams, which plot the expected pixel accuracy as a the function of confidence, and calculate the Expected Calibration Error (ECE) [28]:
@@
where we divide the bins with an interval of 0.1. Note that the reliability diagram is obtained through the bit-wise way in ECOCSeg, meaning that every pixel will provide $K$ samples in total. As observed, ECOCSeg demonstrates smaller gaps between the expected accuracy and confidence, indicating superior calibration of predictions. While the one-hot encoding form is typically notorious for inflating the probability of the predicted class [70] and suffers higher calibration error accordingly, ECOCSeg exhibits better reliability and interpretability through fine-grained bit-level label representation.
@@
$[1]$ [7] Weakly-supervised Discriminative regions yield incomplete pseudo-labels Alternating optimization: 1) cluster sub-classes per class; 2) train a sub-class classifier Sub-class discovery at the feature level
$[2]$ [40] Domain adaptive Prior works focus on intra-class alignment without explicitly modeling inter-class structure Online prototype update for contrastive learning Structure modeling via class prototypes
$[3]$ [100] Domain adaptive Improve pseudo-labels via feature clustering Prototype-based denoising Feature-space method
$[4]$ [71] Domain adaptive Avoid manual thresholds Symmetric distillation & consensus Threshold-free, feature-level
$[5]$ [98] Domain adaptive Fuse online-offline pseudo-labels Unified multi-branch fusion Fusion in feature space
$[6]$ [62] Active Domain adaptive Using a small amount of labeled target data to guide adaptation Anchor-based soft alignment Feature-space alignment
$[7]$ [59] Semi-supervised Treat prediction error as a learnable correction term Introduce a correction network that learns to refine predictions with residual errors. Prediction refinement via residual correction
$[8]$ [24] Semi-supervised Reduce confirmation bias Improve pseudo-labels by peeking at future model states Modification of the teacher model
@@
Figure 10: More qualitative comparison built with UniMatch [93] on SSL benchmark of Pascal. The significant improvements are marked with dotted boxes.